REU 2006: SCHEDULE FOR WEEKS ONE AND TWO

ALL TALKS WILL BE IN ECKHART RY251 (unless posted otherwise)

MONDAY, June 19:

Welcome and opening meeting: 9:00 in RYERSON 251. Pizza and coke: 12:00 in the barn. Tutorial organizational meeting: 1:00 in the barn.

FIRST WEEK:

Date Abert Babai Farb Fefferman May
Mon, June 19 10:30–12:00  9:15–10:15 4:00-5:00
Tue, June 20  9:00–12:00 1:00–3:00
Wed, June 21 10:30–12:00  9:00–10:15 4:00-5:00
Thu, June 22  9:00–12:00 1:00–3:00
Fri, June 23 10:30–12:00  9:00–10:15 4:00-5:00

TUTORIALS: 1:00–4:00 and 5:00–? MWF; 3:00–? TTh

SECOND WEEK:

NOTE: attendance at the YSP and SESAME morning preparation and training sessions by all people who will be counsellors in either of these programs is required.

Date  YSP/SES Abert/Babai  Abert  Babai  Farb  Fefferman  May
Apprentice Problems Discrete
Mon, June 26 10:20–1:00 10:30–12:00 2:00–3:30 9:00–10:15 4:00–5:00
Tue, June 27 9:00–12:00 9:30–12:00 3:30–5:00  1:00–3:00
Wed, June 28 10:20–1:00 10:30–12:00 2:00–3:30 9:00–10:15 4:00–5:00
Thu, June 29 9:00–12:00 9:30–12:00 3:30–5:00  1:00–3:00
Fri, June 30  9:00–12:00 9:30–12:00  2:00–3:30  4:00–5:00

TUTORIALS: 5:00 onwards

Here is the schedule in summary form:

FIRST WEEK:

Fefferman (Analysis and game theory) 9:15–10:15 M and 9:00–10:15 WF Abert (Problems and theorems): 10:30–12:00 MWF Babai (Discrete mathematics): 9:00–12:00 TTh Farb (Geometry and topology): 1:00–3:00 TTh May (K-theory): 4:00–5:00 MWF

SECOND WEEK:

YSP/SESAME Preparation: 10:20–1:00 MW, 9:00–12:00 TThF Fefferman (analysis and game theory): 9:00–10:15 MW Abert/Babai (Apprentice): 10:30–12:00 MW, 9:30-12:00 TThF Abert (Problems): 3:30–5:00 TTh Babai (Discrete mathematics): 2:00–3:30 MWF Farb (Geometry and topology): 1:00–3:00 TTh May (K-theory): 4:00-5:00 MWF